solve each quadratic in form equation

2. x-3x^1/2+2=0
1
Expert's answer
2012-12-19T10:44:09-0500
let t=x^1/2 then we have
t^2-3t+2=0
2 roots t1=1, t2=2
x^1/2=1 or x^1/2=2
x=1 or x=4